# Posterior Ramifications of Prior Dependence Structures

### 11 Dec 2023 v1

@@ 1.3 Contributions
This paper provides general guidance for prior dependence specification in multivariate settings. These recommendations are topical given that recent advances in copula-based priors allow for the incorporation of unprecedented flexibility into the prior dependence structure. While this additional flexibility could give rise to priors that more accurately characterize real-life phenomena, practitioners may be subject to choice overload when deciding between the vast number of potential dependence structures. We argue that this additional flexibility is only appropriate for certain prior distributions. Our recommendations help practitioners discard prior dependence structures that are not suitable for posterior analysis – simplifying the prior specification process. These recommendations also clarify how the prior dependence structure can impact the posterior distribution. Our recommendations are illustrated using several models for which copula-based priors have been proposed. Flexible copula-based methods continue to be proposed as possible extensions in new settings (Perepolkin et al., 2023), so this paper’s commentary is broadly applicable. Unlike situations that require the specification of a community of prior distributions, we restrict our discussion to the case where a single prior is elicited.
The remainder of this article is structured as follows. In Section 2, we define general conditions under which the prior dependence structure is incompatible with that induced by the likelihood function and unable to be retained by the posterior distribution as data are observed. We formally prove this result and demonstrate that priors elicited using the methods proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) are generally incompatible with standard likelihood functions. We clarify the settings for which this additional flexibility in the prior dependence structure is useful in Section 3. In Section 4, we prove several theoretical results about the impact of the prior dependence structure on the convergence of the posterior mode, which we contrast with previous work on copula-based priors (Michimae and Emura, 2022). We then conduct numerical studies to contextualize these theoretical results, which prompt general recommendations regarding choosing a prior dependence structure for posterior analysis. We provide concluding remarks and a discussion of extensions to this work in Section 5.
2 The Prior-Posterior Disconnect
@@ 2.1 Background
In this section, we examine situations where the prior dependence structure cannot carry over into the posterior distribution as data are collected. It is unrealistic to expect the prior dependence structure for $\boldsymbol{\theta}$ to be perfectly specified. However, it presents practical issues if we select a prior dependence structure that cannot be retained a posteriori and plan to use the posterior of $\boldsymbol{\theta}$ to draw conclusions about its dependence structure. This would lead us to draw different conclusions about the dependence structure given small and large samples generated from the same data generation process, giving rise to a disconnect between the prior and posterior. Such prior dependence structures also make it problematic to consider the posterior of a function of several parameters in $\boldsymbol{\theta}$.
@@
We introduce the concept of chronic rejection (Libby and Pober, 2001; Vos et al., 2011) to frame this section’s main result. The term chronic rejection is often used to describe the process in which a transplanted organ is rejected by the recipient’s immune system over a long period of time. The recipient’s persistent immune response against the transplanted organ causes slow and gradual damage. This is in contrast to acute or hyperacute rejection that occurs shortly after the organ is transplanted. Chronic rejection is a more appropriate analogy for this work because we may require very large sample sizes for the vestiges of the prior dependence structure to disappear. Similar to finding a suitable organ donor, we argue that one should consider whether the prior dependence structure for $\boldsymbol{\theta}$ is compatible with that induced by the likelihood function. Such incompatibilities imply that the prior dependence structure cannot be retained a posteriori – and is hence chronically rejected. We emphasize that the term rejection as used in this paper does not imply the formal rejection of a statistical hypothesis test. In Section 2.2, we define broad conditions under which the posterior chronically rejects the prior dependence structure. We illustrate this phenomenon using simulation in Section 2.3.
2.2 Chronic Rejection of Prior Dependence in Posterior Distributions
@@
Here, we define general conditions under which prior dependence structures are chronically rejected by the posterior. These sufficient conditions can be readily verified prior to observing data. When planning posterior analyses, practitioners can discard prior dependence structures that satisfy these conditions. This simplification reduces the number of potential dependence structures and hence streamlines the prior elicitation process.
@@
We consider the limiting behaviour of the posterior distribution for $\boldsymbol{\theta}$. Because we consider this behaviour under various data generation processes, the data are random variables. Data from a random sample of size $n$ are represented by $\boldsymbol{Y}^{{}_{(n)}}$. Realizations of these samples are denoted by $\boldsymbol{y}^{{}_{(n)}}$. Much of the work on limiting posteriors (see e.g., (Ghosal et al., 1995), (Le Cam and Yang, 2000), (Gao et al., 2020), or (Schillings et al., 2020)) appeals to the Bernstein-von Mises theorem (van der Vaart, 1998). We let $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ be the statistical model corresponding to the likelihood function in (1). When data $\boldsymbol{Y}^{{}_{(n)}}$ are generated independently and identically from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$, the Bernstein-von Mises theorem dictates that the posterior for $\boldsymbol{\theta}$ converges to the $\mathcal{N}(\boldsymbol{\theta}_{0},\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}/n)$ distribution in the limit of infinite data, where $\mathcal{I}(\cdot)$ is the Fisher information.
@@
$p(\boldsymbol{\theta})=\prod_{j=1}^{d}f_{j}(\theta_{j})\times c(F_{1}(\theta_{1}),...,F_{d}(\theta_{d})),$ (4)
where $c(u_{1},...,u_{d})$ is the copula density function for an absolutely continuous copula $C$ and $F_{j}(\theta_{j})$ and $f_{j}(\theta_{j})$ are respectively the marginal prior CDF and PDF for $\theta_{j}$, $j=1,...,d$. Our first result in Corollary 1 is a direct consequence of the Bernstien-von Mises theorem.
Corollary 1.
Let the model $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$ and prior $p(\boldsymbol{\theta})$ satisfy all conditions for the Bernstien-von Mises theorem. Let $\boldsymbol{Y}^{{}_{(n)}}$ be generated independently from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$. The posterior dependence structure of $\boldsymbol{\theta}$ given observed $\boldsymbol{y}^{{}_{(n)}}$ converges to $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ as $n\rightarrow\infty$.
By the Bernstein-von Mises theorem, the posterior of $\boldsymbol{\theta}$ is approximately multivariate normal with covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}/n$, so its dependence structure is reasonably characterized by a Gaussian copula with correlation matrix $\boldsymbol{R}$ corresponding to $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$. We next extend this result to a more general setting where the observations in $\boldsymbol{Y}^{{}_{(n)}}$ are independently but not identically distributed. We now explore the limiting behaviour of the posterior when data are generated from the following prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$:
[…]
where $p_{D}(\boldsymbol{\theta})$ is a prior that characterizes the data generation process for $\boldsymbol{Y}^{{}_{(n)}}$. The prior in (5) may or may not be the same prior as $p(\boldsymbol{\theta})$ used to analyze the observed data in (4), which is often called the analysis prior. The Bernstein-von Mises theorem considers a special case of $p(\boldsymbol{y}^{{}_{(n)}})$ in which the prior $p_{D}(\boldsymbol{\theta})$ is degenerate. That is, $p_{D}(\boldsymbol{\theta}_{0})=1$ for a particular $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}$, and 0 otherwise. In light of this, we emphasize that the prior that must satisfy condition 3 for the Bernstein-von Mises theorem is the analysis prior $p(\boldsymbol{\theta})$. Theorem 1 generalizes the result from Corollary 1 for nondegenerate priors $p_{D}(\boldsymbol{\theta})$ under certain conditions.
@@
Let $\boldsymbol{\Theta}^{*}$ be the set of interior points in $\boldsymbol{\Theta}$. Suppose conditions 1, 2, and 3 for the Bernstein-von Mises theorem are satisfied for all possible $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}^{*}$. Let the likelihood function $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y})$ be globally log-concave for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}^{*}$. Let data $\boldsymbol{Y}^{{}_{(n)}}$ be generated independently from $p(\boldsymbol{y}^{{}_{(n)}})$ in (5) where $p_{D}(\boldsymbol{\theta})=0$ for all $\boldsymbol{\theta}\notin\boldsymbol{\Theta}^{*}$. For sufficiently large $n$, the posterior dependence structure of $\boldsymbol{\theta}$ given observed $\boldsymbol{y}^{{}_{(n)}}$ is approximately characterized by $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}$ for the posterior mode $\tilde{\boldsymbol{\theta}}=\text{arg\,max}_{\boldsymbol{\theta}}~{}p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$.
Proof of Theorem 1. The degeneracy of $p_{D}(\boldsymbol{\theta})$ is the only condition of the Bernstein-von Mises theorem that Theorem 1 relaxes: each observation can be generated independently from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ parameterized with a different value for $\boldsymbol{\theta}\sim p_{D}(\boldsymbol{\theta})$. The first sentence in Theorem 1 ensures that all other conditions for the Bernstein-von Mises theorem are satisfied for all interior points in $\boldsymbol{\Theta}$. Unlike in Corollary 1, we now require that the likelihood function $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ be globally log-concave for all $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}^{*}$, which ensures that the posterior of $\boldsymbol{\theta}$ is unimodal with mode $\tilde{\boldsymbol{\theta}}$ given large samples $\boldsymbol{y}^{{}_{(n)}}$. When the conditions for Theorem 1 are satisfied, $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ is approximated by the following multivariate normal distribution (Gelman et al., 2013) for sufficiently large $n$:
$\mathcal{N}\left(\tilde{\boldsymbol{\theta}},\left[-\dfrac{\partial^{2}}{\partial\boldsymbol{\theta}^{2}}\text{log}(p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))\bigg{\lvert}_{\boldsymbol{\theta}=\tilde{\boldsymbol{\theta}}}\right]^{-1}\right).$ (6)
The likelihood function’s impact on the posterior of $\boldsymbol{\theta}$ overwhelms that of the prior as the sample size $n$ increases. For sufficiently large $n$, we can therefore replace $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ in the variance term of (6) with $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$. It follows that the posterior of $\boldsymbol{\theta}$ given observed $\boldsymbol{y}^{{}_{(n)}}$ is approximately $\mathcal{N}(\tilde{\boldsymbol{\theta}},\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}/n)$ for sufficiently large samples $\boldsymbol{y}^{{}_{(n)}}$. Thus, the posterior dependence structure is approximately characterized by the Gaussian copula corresponding to the covariance matrix $\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}$.
@@
Suppose the conditions for Theorem 1 are satisfied. Let $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ describe prior dependence on $\boldsymbol{\theta}$. If no $\boldsymbol{\theta}\in\boldsymbol{\Theta}^{*}$ is such that $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta})^{-1}$ induces a dependence structure $\{\tau(\boldsymbol{\theta})_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ such that $\tau(\boldsymbol{\theta})_{e_{1},e_{2}|D_{\boldsymbol{e}}}=\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}$ for all $\boldsymbol{e}\in E(\mathcal{V})$, the posterior of $\boldsymbol{\theta}$ cannot retain the magnitude and direction of prior dependence as data $\boldsymbol{Y}^{{}_{(n)}}$ are generated independently from $p(\boldsymbol{y}^{{}_{(n)}})$ in (5) where $p_{D}(\boldsymbol{\theta})=0$ for all $\boldsymbol{\theta}\notin\boldsymbol{\Theta}^{*}$.
@@
Corollary 2 follows directly from Theorem 1. We suppose that there is no $\boldsymbol{\theta}\in\boldsymbol{\Theta}^{*}$ such that the Gaussian copula corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta})^{-1}$ induces a dependence structure characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$. Thus, it is impossible for the Gaussian copula corresponding to the covariance matrix $\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}$ to induce a dependence structure with that partial characterization for $\tilde{\boldsymbol{\theta}}\in\boldsymbol{\Theta}^{*}$. It follows that the magnitude and direction of the dependence structure for $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ cannot be characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ for sufficiently large $n$. We emphasize that considering dependence structures via Kendall’s tau on the D-vine structure of the copula does not fully specify the dependence structure. This allows for more flexibility in the choices for the copula families; it also facilitates the consideration of prior dependence for subvectors of $\boldsymbol{\theta}$, which is useful because it may require too much cognitive complexity to assess the full prior dependence structure. However, two bivariate copulas may have the same Kendall’s tau but different properties in terms of symmetry and tail dependence. We focus on the magnitude and direction of dependence to present generally applicable guidance for prior specification. Corollary 2 therefore provides a straightforward result that can be used to discard potential dependence structures for $\boldsymbol{\theta}$ that will be chronically rejected as data $\boldsymbol{y}^{{}_{(n)}}$ are observed.
We now illustrate how Corollary 2 can be applied in several scenarios. We first consider prior specification for $\boldsymbol{\theta}=(\mu,\sigma^{2})$ in the typical $\mathcal{N}(\mu,\sigma^{2})$ model. For this model, the inverse Fisher information matrix $\mathcal{I}(\boldsymbol{\theta})^{-1}$ is diagonal for all possible values of $\mu\in\mathbb{R}$ and $\sigma^{2}\in\mathbb{R}^{+}$. It follows by Corollary 2 that the joint posterior of $\mu$ and $\sigma^{2}$ cannot retain a prior dependence structure for which Kendall’s tau is not 0. Under the conditions for Theorem 1, any attempt to inject positive or negative dependence into this posterior will be unsuccessful in the limit of infinite data. We also consider the standard gamma model parameterized by $\boldsymbol{\theta}=(\alpha,\beta)$, where $\alpha$ and $\beta$ are respectively the shape and rate parameters. The correlation between $\alpha$ and $\beta$ dictated by the inverse Fisher information matrix is $1/\sqrt{\alpha\psi_{1}(\alpha)}$, where $\psi_{1}(\cdot)$ is the trigamma function. The correlation $1/\sqrt{\alpha\psi_{1}(\alpha)}$ is a positive and increasing function for all $\alpha>0$. When the conditions for Theorem 1 hold, the joint posterior of $\boldsymbol{\theta}$ will be unable to retain negative dependence structures between $\alpha$ and $\beta$. The last model considered in this subsection is the multinomial model. In Section SM1 of the supplementary materials, we show that the inverse Fisher information matrix for the multinomial model parameterized in terms of $\boldsymbol{\theta}=(Z_{1},Z_{2},...,Z_{w-1})$ from (2) is diagonal for all possible $(Z_{1},Z_{2},...,Z_{w-1})\in(0,1)^{w-1}$. When combined with the standard multinomial likelihood, the prior specification methods proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) satisfy the conditions for Theorem 1. If $p(\boldsymbol{\theta})$ incorporates any positive or negative dependence between the parameters, this dependence structure will be rejected as multinomial data $\boldsymbol{y}^{{}_{(n)}}$ are observed – even when $p_{D}(\boldsymbol{\theta})$ from the prior predictive distribution in (5) is the same as the analysis prior $p(\boldsymbol{\theta})$!

### 23 Mar 2024 v2

@@ 1.4 Contributions
This paper provides general guidance for prior dependence specification in multivariate settings. These recommendations are topical given that recent advances in copula-based priors allow for the incorporation of unprecedented flexibility into the prior dependence structure. While this additional flexibility could give rise to priors that more accurately characterize real-life phenomena, practitioners may be subject to choice overload when deciding between the vast number of potential dependence structures. We argue that this additional flexibility is only useful in certain contexts. Our recommendations help practitioners discard prior dependence structures that do not align with their objectives for posterior analysis – simplifying the prior specification process. These recommendations also clarify how the prior dependence structure can impact the posterior distribution. Our recommendations are illustrated using several models for which copula-based priors have been proposed. Unlike situations that require the specification of a community of prior distributions, we restrict our discussion to the case where a single prior is elicited.
The remainder of this article is structured as follows. In Section 2, we define general conditions under which the prior dependence structure is incompatible with that induced by the likelihood function and unable to be retained by the posterior distribution as data are observed. We formally prove this large-sample result and demonstrate that priors elicited using the methods proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) are generally incompatible with standard likelihood functions. We discuss small-sample scenarios where this additional flexibility in the prior dependence structure is nevertheless useful in Section 3. In Section 4, we prove asymptotic results about the impact of the prior dependence structure on the convergence of the posterior mode, which we contrast with previous work on copula-based priors (Michimae and Emura, 2022). We then conduct numerical studies with both small and large sample sizes to contextualize these theoretical results. These simulations prompt further recommendations regarding choosing a prior dependence structure for posterior analysis. We provide concluding remarks and a discussion of extensions to this work in Section 5.
2 Retention of Prior Dependence
@@ 2.1 Background
In this section, we examine situations where the prior dependence structure cannot carry over into the posterior distribution as data are collected. It is unrealistic to expect the prior dependence structure for $\boldsymbol{\theta}$ to be perfectly specified. However, we should be mindful of whether the prior dependence structure can be retained a posteriori before using the posterior of $\boldsymbol{\theta}$ to draw conclusions about its dependence structure. Otherwise, we may draw different conclusions about the dependence structure given small and large samples generated from the same data generation process. Drawing such conclusions in an uninformed way presents practical issues and complicates the consideration of the posterior of a function of several parameters in $\boldsymbol{\theta}$.
@@
The concept of chronic rejection (Libby and Pober, 2001; Vos et al., 2011) frames this section’s main result. The term chronic rejection describes the process in which a transplanted organ is rejected by the recipient’s immune system over a long period of time. The recipient’s persistent immune response against the transplanted organ causes gradual damage. Similar to screening an organ donor, we argue that one should consider whether the prior dependence structure for $\boldsymbol{\theta}$ is compatible with that induced by the likelihood function. Such incompatibilities imply that the prior dependence structure cannot be retained a posteriori – and the prior dependence structure is hence a chronically rejected one. We emphasize that the term rejection as used in this paper does not imply the formal rejection of a statistical hypothesis test. In Section 2.2, we define the notion of chronically rejected prior dependence structures. We illustrate this definition using simulation in Section 2.3. Section 3 discusses situations where using chronically rejected prior dependence structures may be sensible, so practitioners should consider their posterior objectives before choosing a prior dependence structure.
2.2 Chronically Rejected Prior Dependence Structures
@@
Here, we define general conditions under which prior dependence structures cannot be retained by the posterior with enough data. These sufficient conditions can be readily verified prior to observing data. When planning posterior analyses, practitioners can discard prior dependence structures that satisfy these conditions for certain posterior objectives. This simplification reduces the number of potential dependence structures and hence streamlines the prior elicitation process.
@@
We consider the limiting behaviour of the posterior distribution for $\boldsymbol{\theta}$. Because we consider this behaviour under various data generation processes, the data are random variables. Data from a random sample of size $n$ are represented by $\boldsymbol{Y}^{{}_{(n)}}$. Realizations of these samples are denoted by $\boldsymbol{y}^{{}_{(n)}}$. Much of the work on limiting posteriors (see e.g., Ghosal et al. (1995), Le Cam and Yang (2000), Gao et al. (2020), or Schillings et al. (2020)) appeals to the Bernstein-von Mises theorem (van der Vaart, 1998). We let $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ be the statistical model corresponding to the likelihood function in (1). When data $\boldsymbol{Y}^{{}_{(n)}}$ are generated independently and identically from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$, the Bernstein-von Mises theorem dictates that the posterior for $\boldsymbol{\theta}$ converges to the $\mathcal{N}(\boldsymbol{\theta}_{0},\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}/n)$ distribution in the limit of infinite data, where $\mathcal{I}(\cdot)$ is the Fisher information.
@@
$p(\boldsymbol{\theta})=\prod_{j=1}^{d}f_{j}(\theta_{j})\times c(F_{1}(\theta_{1}),\dots,F_{d}(\theta_{d})),$ (4)
where $c(u_{1},\dots,u_{d})$ is the copula density function for an absolutely continuous copula $C$ and $F_{j}(\theta_{j})$ and $f_{j}(\theta_{j})$ are respectively the marginal prior CDF and PDF for $\theta_{j}$, $j=1,\dots,d$.
To directly apply the Bernstein-von Mises theorem, a single value of $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}$ must be selected. It is not realistic to expect practitioners to correctly identify this value for $\boldsymbol{\theta}_{0}$ a priori. As such, our results incorporate uncertainty about the value of $\boldsymbol{\theta}_{0}$ used to generate $\boldsymbol{Y}^{{}_{(n)}}$. We do so by introducing a prior $p_{D}(\boldsymbol{\theta})$ that defines the prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$:
$p(\boldsymbol{y}^{{}_{(n)}})=\int\prod_{i}^{n}m(y_{i}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})\hskip 1.0ptp_{D}(\boldsymbol{\theta})\hskip 1.0ptd\boldsymbol{\theta}.$ (5)
In (5), data $\boldsymbol{Y}^{{}_{(n)}}$ are generated independently and identically from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$ given $\boldsymbol{\theta}_{0}\sim p_{D}(\boldsymbol{\theta})$. This data generation process allows us to compare various objectives for prior specification using repeated simulation in Section 3. The prior in (5) may or may not be the same prior as $p(\boldsymbol{\theta})$ used to analyze the observed data in (4), which is often called the analysis prior. To explore the limiting behaviour of the posterior when data are generated via (5), we note that the Bernstein-von Mises theorem considers a special case of $p(\boldsymbol{y}^{{}_{(n)}})$ in which the prior $p_{D}(\boldsymbol{\theta})$ is degenerate. That is, $p_{D}(\boldsymbol{\theta}_{0})=1$ for a particular $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}$, and 0 otherwise. In light of this, we emphasize that the prior that must satisfy condition 3 for the Bernstein-von Mises theorem is the analysis prior $p(\boldsymbol{\theta})$. Theorem 1 generalizes results from the Bernstein-von Mises theorem to nondegenerate priors $p_{D}(\boldsymbol{\theta})$ under certain conditions.
@@
Let $\boldsymbol{\Theta}^{*}$ be the set of interior points in $\boldsymbol{\Theta}$. Suppose conditions 1, 2, and 3 for the Bernstein-von Mises theorem are satisfied for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}^{*}$. Let data $\boldsymbol{Y}^{{}_{(n)}}$ be generated via (5) such that $\boldsymbol{\theta}_{0}\sim p_{D}(\boldsymbol{\theta})$ and $p_{D}(\boldsymbol{\theta})=0$ for all $\boldsymbol{\theta}\notin\boldsymbol{\Theta}^{*}$. The posterior dependence structure of $\boldsymbol{\theta}$ given observed $\boldsymbol{y}^{{}_{(n)}}$ converges to $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ as $n\rightarrow\infty$.
Proof of Theorem 1. When the conditions for Theorem 1 hold, $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ is approximately multivariate normal with covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}/n$ for large enough $n$. The posterior dependence structure of $\boldsymbol{\theta}$ is hence reasonably characterized by a Gaussian copula with correlation matrix $\boldsymbol{R}$ corresponding to $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$. This result follows by the Bernstein-von Mises theorem because we restrict the value of $\boldsymbol{\theta}_{0}$ to be contained in $\boldsymbol{\Theta}^{*}$; the Bernstein-von Mises theorem is applicable when $\boldsymbol{\theta}_{0}$ is not a boundary point of the parameter space $\boldsymbol{\Theta}$.
@@
Under the conditions for Theorem 1, let $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ describe prior dependence on $\boldsymbol{\theta}$. Suppose no $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ is such that $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ induces a dependence structure $\{\tau(\boldsymbol{\theta})_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ such that $\tau(\boldsymbol{\theta}_{0})_{e_{1},e_{2}|D_{\boldsymbol{e}}}=\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}$ for all $\boldsymbol{e}\in E(\mathcal{V})$. Then, the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{Y}^{{}_{(n)}}$ cannot retain the magnitude and direction of prior dependence as $n\rightarrow\infty$.
@@
Corollary 1 follows directly from Theorem 1. We suppose there is no $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ such that the Gaussian copula corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ induces a dependence structure characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$. It follows that the magnitude and direction of the dependence structure for $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ cannot be characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ for sufficiently large $n$. We emphasize that considering dependence structures via Kendall’s $\tau$ on the D-vine structure of the copula does not fully specify the dependence structure. This allows for more flexibility in the choices for the copula families; it also facilitates the consideration of prior dependence for subvectors of $\boldsymbol{\theta}$, which is useful because it may require too much cognitive complexity to assess the full prior dependence structure. However, two bivariate copulas may have the same Kendall’s $\tau$ but different properties in terms of symmetry and tail dependence. We focus on the magnitude and direction of dependence to present generally applicable guidance for prior specification.
Corollary 1 therefore provides a straightforward result that can be used to discard potential dependence structures for $\boldsymbol{\theta}$ that cannot be retained as data $\boldsymbol{y}^{{}_{(n)}}$ are observed. We refer to prior dependence structures that satisfy the conditions for Corollary 1 as chronically rejected dependence structures. The conditions for Corollary 1 are sufficient in that there is no value for $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ such that these dependence structures will be retained as data are generated via (5). However, it is not guaranteed that the prior dependence structure will be retained when the conditions for Corollary 1 are not satisfied. That is, the true value of value $\boldsymbol{\theta}_{0}$ might not be one of those that prevent the conditions for Corollary 1 from being satisfied. Thus, a prior dependence structure that is not retained is not necessarily a chronically rejected one. In this paper, the notion of chronic rejection defines a specific class of prior dependence structures, whereas retention of the prior dependence structure is viewed more generally as a posterior outcome.